Please help me solve this word problem equation,
A store mixes Kenyan coffee worth $10 per kilogram and Turkish coffee worth $12 per kilogram.The mixture is to sell for $11 per kilogram.Find how much of each should be used to make a 588-kilogram mixture.
How many kilograms of the Kenyan coffee should be in the mixture?
How many kilograms of the Turkish coffee should be in the mixture?

Answer by checkley77(12844)

12X+10(588-X)=588*11
12X+5,880-10X=6,468
2X=6,468-5,880
2X=588
X=588/2
X=294 POUNDS OF $12 COFFEE IS USED.
588-294=294 LOUNDS OF $10 COFFEE IS USED.
PROOF:
12*294+10*294=588*11
3,528+2,940=6,468
6,468=6,468